355 people used the public swimming pool. The daily prices are $1.50 for children and $2.50 for adults. The receipts for admission totaled .$646.50. How many children and how many adults swam at the public pool that​ day?
Answer by jorel555(1290)

Let n be the number of adults. Then
2.5n+1.5(355-n)=646.5
2.5n+532.5-1.5n=646.5
n=646.5-532.5=114
There were 114 adults, and 241 children that day. ☺☺☺☺